# Scaling Off-Policy Reinforcement Learning with Batch and Weight Normalization

### v1

@@ 4 Combining batch normalization and weight normalization for scaling up
Theorem 4.2 (Van Laarhoven (2017)).
Let $f(\bm{X};\bm{w},b,\gamma,\beta)$ be a scale-invariant function. Then, the gradients of $f$ scale inversely proportional to the scaling factor $c\in\mathbb{R}$ of its parameters $\bm{w}$,
$\nabla f(\bm{X};c\bm{w},cb,\gamma,\beta)=\nabla f(\bm{X};\bm{w},b,\gamma,\beta)/c.$ (6)
Proof. Appendix B
Recently, Lyle et al. (2024) demonstrated that the combination of ln and wn can help mitigate loss of plasticity. Since the gradient scale is inversely proportional to $c$, keeping norms constant helps to maintain a stable effective learning rate (elr,Van Hasselt et al. (2019)), further enhancing training stability.
We conjecture that maintaining a stable elr could also be beneficial when increasing the utd ratios in continuous control rl. As the utd ratio increases, the networks are updated more frequently with each environment interaction. Empirically, we find that the network norms tend to grow quicker with increased utd ratios (Figure 2), which in turn decreases the elr even quicker and could be the case for instabilities and low training performance. As such, we empirically investigate the effectiveness of combining CrossQ with wn with increasing utd ratios.
@@
We apply wn to the first two linear layers, ensuring that their weights remain unit norm after each gradient step by projecting them onto the unit ball, similar to Lyle et al. (2024). Since this constraint effectively stabilizes the elr in these layers, we found it beneficial to slightly reduce the learning rate from $1$e33-3- 3 to $1$e44-4- 4. While we could employ a learning rate schedule (Lyle et al., 2024) we did not investigate this here. Additionally, we impose weight decay on all parameters that remain unbounded—specifically the final dense output layer. In practice, we use AdamW (Loshchilov, 2017) with a decay of 00 (which falls back to vanilla Adam (Kingma, 2014)) for the normalized intermediate dense layers and $1$e22-2- 2 otherwise. We chose a maximum utd of $5$ for our experiments due to our computational budget and good strong in sample efficiency.
@@
CrossQ famously removed the target networks from the actor-critic framework and showed that using bn training remains stable even without them (Bhatt et al., 2024). While we find this to be true in many cases, we find that especially in dmc, the re-integration of target networks can help stabilize training overall (see Section 5.4). However, not surprisingly, we find that the integration of target networks with bn requires careful consideration of the different state-action distributions between the $\bm{s},\bm{a}$ and $\bm{s}^{\prime},\bm{a}^{\prime}\sim\pi(\bm{s}^{\prime})$ exactly as proposed by Bhatt et al. (2024). To satisfy this, we keep the joined forward pass through both the critic network as well as the target critic network. We evaluate both networks in training mode, i.e., they calculate the joined state-action batch statistics on the current batches. As is common, we use Polyak-averaging with a $\tau=0.005$ from the critic network to the target network.
5 Experiments
To evaluate the effectiveness of our proposed CrossQ + wn method, we conduct a comprehensive set of experiments on the DeepMind Control Suite (Tassa et al., 2018) and MyoSuite (Caggiano et al., 2022) benchmarks. Our primary goal is to investigate the scalability of CrossQ + wn with increasing utd ratios and to assess the stabilizing effects of combining CrossQ with wn. We compare our approach to several baselines, including the recent bro (Nauman et al., 2024), CrossQ (Bhatt et al., 2024) and sr-sac (D’Oro et al., 2022) a version of sac (Haarnoja et al., 2018) with high utd ratios and network resets.
@@ 5.1 Experimental setup
Our implementation is based on the sac implementation of jaxrl codebase (Kostrikov, 2021). We implement CrossQ following the author’s original codebase and add the architectural modifications introduced by (Bhatt et al., 2024), incorporating batch normalization in the actor and critic networks. We extend this approach by introducing wn to regulate the growth of weight norms and prevent loss of plasticity and add target networks. We perform a grid search to focus on learning rate selection and layer width.
We evaluate $25$ diverse continuous control tasks, $15$ from dmc and $10$ from MyoSuite. These tasks vary significantly in complexity, requiring different levels of fine motor control and policy adaptation with high dimensional state spaces up to $\mathbb{R}^{223}$.
Each experiment is run for 1 million environment steps and across $10$ random seeds to ensure statistical robustness. We evaluate agents every $25,000$ environment steps for $5$ trajectories. As proposed by Agarwal et al. (2021), we report the interquartile mean (iqm) and $95\%$ stratified bootstrap confidence intervals (cis) of the return, if not otherwise stated.
For the bro baseline results, for computational reasons, we take the official evaluation data the authors provide. The official bro codebase is also based on jaxrl, and the authors followed the same evaluation protocol, making it a fair comparison.
Figure 3: CrossQ wn + utd$\mathbin{=}5$ against baselines. We compare our proposed CrossQ + wn utd$\mathbin{=}5$ against two baselines, bro (Nauman et al., 2024) and sr-sac utd$\mathbin{=}32$. Results are reported on all $15$ dmc and $10$ Myosuite tasks. We plot the iqm and $95\%$ cis over $10$ random random seeds. Our proposed approach proves competitive to bro and outperforms the CrossQ baseline. We want to note that our approach achieves this performance without requiring any parameter resetting or additional exploration policies.
5.2 Weight normalization allows CrossQ to scale effectively
We provide empirical evidence for our hypothesis that controlling the weight norm and, thereby, the elr can stabilize training. We show that through the addition of wn, CrossQ + wn shows stable training and can stably scale with increasing utd ratios.
@@ 5.2 Weight normalization allows CrossQ to scale effectively
These results show that CrossQ + wn utd$\mathbin{=}5$ is competitive to the bro baseline on both dmc and Myosuite, especially on the more complex dog and humanoid tasks. Notably, CrossQ + wn utd$\mathbin{=}5$ uses only half the utd of bro and does not require any parameter resets and no additional exploration policy. Further, it uses $\sim 90\%$ fewer network parameters—bro reports $\sim 5M$, while our proposed CrossQ + wn uses only $\sim 600k$ (these numbers vary slightly per environment, depending on the state and action dimensionalities).
In contrast, vanilla CrossQ utd$\mathbin{=}1$ exhibits much slower learning on most tasks and, in some environments, fails to learn performant policies. Moreover, the instability of vanilla CrossQ at utd$\mathbin{=}5$ is particularly notable, as it does not reliably converge across environments.
@@ 5.3 Stable scaling of CrossQ + wn with utd ratios
Figure 4: CrossQ wn utd scaling behavior. We plot the iqm return and $95\%$ confidence intervals for different utd ratios $\in\{1,2,5\}$. The results are aggregated over 15 dmc environments and $10$ random seeds each according to Agarwal et al. (2021). The sample efficiency scales reliably with increasing utd ratios.
To visualize the stable scaling behavior of CrossQ + wn we ablate across three different utd ratios $\in\{1,2,5\}$. Figure 4 shows training curves aggregated over all $15$ dmc tasks. As expected, CrossQ + wn shows reliable scaling behavior, with the learning curves ordered in increasing order accordance to their respective utd ratio.
@@ 5.4 Hyperparameter ablation studies
We also ablate the different hyperparameters of CrossQ + wn utd$\mathbin{=}5$, by changing each one at a time. Figure 5 shows aggregated results of the final performances of each ablation. We will briefly discuss each ablation individually.
@@
Not performing weight normalization results in the biggest drop in performance across all our ablations. This loss is most drastic on the Myosuite tasks and often results in no meaningful learning. Showing that, as hypothesized, the inclusion of wn into the CrossQ framework yields great improvements in terms of sample efficiency and training stability, especially for larger utd ratios.
Update-to-data ratio.
As expected, decreasing the utd ratio decreases the performance of CrossQ + wn, as demonstrated in the previous section. Aggregated over all dmc environments, this effect is the smallest, since this aggregation includes easier environments as well. Looking at the harder dog and humanoid tasks, as well as Myosuite, the effect is more pronounced. However, lower utds are already reasonably competitive in overall performance.
@@
Ablating the target networks shows that on Myosuite, there is no significant difference between using a target network and or no target network. Results on dmc differ significantly. There, removing target networks leads to a significant drop in performance, nearly as large as removing weight normalization. This finding is interesting, as it suggests that CrossQ + wn without target networks is not inherently unstable. But there are situations where the inclusion of target networks is required. Further investigating the role and necessity of target networks in rl is an interesting direction for future research.
@@ 6 Related work
Model-based rl
methods enhance sample efficiency by constructing predictive models of the environment to reduce reliance on real data collection (Sutton, 1990; Janner et al., 2019; Feinberg et al., 2018; Heess et al., 2015). However, such methods introduce additional complexity, computational overhead, and potential biases due to model inaccuracies.
@@
Model-free rl methods, including those utilizing higher utd ratios, focus on increasing the number of gradient updates per collected sample to maximize learning from available data. High utd training introduces several challenges, such as overfitting to early training data, a phenomenon known as primacy bias (Nikishin et al., 2022). This can be counteracted by periodically resetting the network parameters (Nikishin et al., 2022; D’Oro et al., 2022). However, network resets introduce abrupt performance drops. Alternative approaches use techniques such as Q-function ensembles (Chen et al., 2021; Hiraoka et al., 2021) and architectural changes (Nauman et al., 2024).
@@
Normalization techniques have long been recognized for their impact on neural network training. ln (Ba et al., 2016) and other architectural modifications have been used to stabilize learning in rl (Hiraoka et al., 2021; Nauman et al., 2024). Yet bn has only recently been successfully applied in this context (Bhatt et al., 2024), challenging previous findings, where bn in critics caused training to fail (Hiraoka et al., 2021). wn has been shown to keep elrs stable and prevent loss of plasticity (Lyle et al., 2024), when combined with ln, making it a promising candidate for integration into existing rl frameworks.
Figure 5: Hyperparameter ablations. We ablate CrossQ + wn with respect to the wn, target networks and different utd.
7 Limitations & future work
In this work, we only consider continuous state-action benchmarking tasks. While our proposed CrossQ + wn performs competitively on these tasks, its performance on discrete state-action spaces or vision-based tasks remains unexplored. We plan to investigate this in future work.
8 Conclusion
@@
In this work, we have addressed the instability and scalability limitations of CrossQ in rl by integrating wn. Our empirical results demonstrate that wn effectively stabilizes training and allows CrossQ to scale reliably with higher utd ratios. The proposed CrossQ + wn approach achieves competitive or superior performance compared to state-of-the-art baselines across a diverse set of $25$ complex continuous control tasks from the dmc and Myosuite benchmarks. These tasks include complex and high-dimensional humanoid and dog environments. This extension preserves simplicity while enhancing robustness and scalability by eliminating the need for drastic interventions such as network resets.
Acknowledgments
@@
We wish to thank Joe Watson for helpful discussions and advice during the project. We would also like to thank Tim Schneider, Cristiana de Farias, João Carvalho and Theo Gruner for proofreading and constructive criticism on the manuscript. This research was funded by the research cluster “Third Wave of AI”, funded by the excellence program of the Hessian Ministry of Higher Education, Science, Research and the Arts, hessian.AI.

### v2

@@ 4 Combining batch normalization and weight normalization enables scaling
Theorem 2 (Van Laarhoven [43])
Let $f(\bm{X};\bm{w},b,\gamma,\beta)$ be a scale-invariant function. Its gradient
$\nabla f(\bm{X};c\bm{w},cb,\gamma,\beta)=\nabla f(\bm{X};\bm{w},b,\gamma,\beta)/c,$ (4)
scales inversely proportional to the scaling factor $c\in\mathbb{R}$ of its parameters $\bm{w}$.
The proof is provided in Appendix B.
Recently, Lyle et al. [29] demonstrated that the combination of ln and wn can help mitigate loss of plasticity. Since the gradient scale is inversely proportional to $c$, keeping norms constant helps to maintain a stable effective learning rate (elr,Van Hasselt et al. [42]), further enhancing training stability. We conjecture that maintaining a stable elr could also be beneficial when increasing the utd ratios in continuous control rl. As the utd ratio increases, the networks are updated more frequently with each environment interaction. Empirically, we find that the network norms tend to grow quicker with increased utd ratios (Figure 3), which in turn decreases the elr even quicker and could be the case for instabilities and low training performance. From this observation, we hypothesize that the training phenomena that affect plasticity also appear when attempting sample-efficient learning with higher utds. This hypothesis suggests that regularization techniques for plasticity could also be used to achieve more sample-efficient RL. As a result, we empirically investigate the effectiveness of combining CrossQ with wn with increasing utd ratios.
@@
We apply wn to the first two linear layers, ensuring that their weights remain unit norm after each gradient step by projecting them onto the unit ball, similar to Lyle et al. [29]. While we could employ a learning rate schedule [29] we did not investigate this here as this would add additional complexity. Additionally, we impose weight decay on all parameters that remain unbounded—specifically, the final dense output layer. In practice, we use AdamW [28] with a decay of 00 (which falls back to vanilla Adam [23]) for the normalized intermediate dense layers and $1$e22-2- 2 otherwise.
@@
CrossQ removes the target networks from the actor-critic framework and showed that using bn training remains stable even without them [4]. While we find this to be true in many cases, we find that especially in dmc, the re-integration of target networks can help stabilize training overall (see Section 5.4). However, not surprisingly, we find that the integration of target networks with bn requires careful consideration of the different state-action distributions between the $\bm{s},\bm{a}$ and $\bm{s}^{\prime},\bm{a}^{\prime}\sim\pi(\bm{s}^{\prime})$ exactly as proposed by Bhatt et al. [4]. To satisfy this, we keep the joined forward pass through both the critic network as well as the target critic network. We evaluate both networks in training mode, i.e., they calculate the joined state-action batch statistics on the current batches. As is common, we use Polyak-averaging with a $\tau=0.005$ from the critic network to the target network.
5 Experimental results
To evaluate the effectiveness of our proposed CrossQ + wn method, we conduct a comprehensive set of experiments on the DeepMind Control Suite [40] and MyoSuite [7] benchmarks. For dmc we report individual results for the hard (dog and humanoid), as well as Medium+Easy (cheetah-run, walker-run, hopper-stand, finger-turn-hard, quadruped-run, fish-swim, hopper-hop, pendulum-swingup) due to their varying difficulties. Equally, we split MyoSuite hard and easy environments. Our primary goal is to investigate the scalability of CrossQ + wn with increasing utd ratios and to assess the stabilizing effects of combining CrossQ with wn. We compare our approach to several baselines, including the recent bro [31], CrossQ [4], td-mpc2 [15], and simba [26], a concurrent approach utilizing layer norm. Figure 6 in the appendix further provides a sr-sac [9] baseline, a version of sac [13] with high utd ratios and network resets.
@@ 5.1 Experimental setup
Our implementation is based on the sac implementation of jaxrl codebase [24]. We implement CrossQ following the author’s original codebase and add the architectural modifications introduced by [4], incorporating batch normalization in the actor and critic networks. We extend this approach by introducing wn to regulate the growth of weight norms and prevent loss of plasticity and add target networks. We perform a grid search to focus on learning rate selection and layer width.
We evaluate $25$ diverse continuous control tasks, $15$ from dmc and $10$ from MyoSuite. These tasks vary significantly in complexity, requiring different levels of fine motor control and policy adaptation with high-dimensional state spaces up to $\mathbb{R}^{223}$. Each experiment is run for 1 million environment steps and across $10$ random seeds to ensure statistical robustness. We evaluate agents every $25,000$ environment steps for $5$ trajectories. For the dmc Medium&Easy and Myo Easy we plot the first $200k$ steps, as all methods learn much faster than the 1 million steps. As proposed by Agarwal et al. [1], we report the interquartile mean (iqm) and $90\%$ stratified bootstrap confidence intervals (cis) of the return (or success rate, respectively), if not otherwise stated. For the bro and simba baseline results, for computational reasons, we take the official evaluation data that the authors provide. The official bro codebase is also based on jaxrl, and the authors followed the same evaluation protocol. All experiments were run on a compute cluster with RTX 3090 and A5000 GPUs, where all $10$ seeds run in parallel on a single GPU via jax.vmap.
5.2 Weight normalization allows CrossQ to scale to harder tasks effectively
We provide empirical evidence for our hypothesis that controlling the weight norm and, thereby, the elr can stabilize training (Figure 3). We show that through the addition of wn, CrossQ + wn shows stable training and can stably scale with increasing utd ratios.
@@ 5.2 Weight normalization allows CrossQ to scale to harder tasks effectively
Scaling results in Figure 4 show that CrossQ + wn is competitive to the bro and simba baselines on both dmc and MyoSuite, especially on the more complex dog and humanoid tasks already on lower utd ratios. Notably, CrossQ + wn utd$\mathbin{=}5$ uses only half the utd of bro and does not require any parameter resets and no additional exploration policy. Further, it uses $\sim 90\%$ fewer network parameters—bro reports $\sim 5M$, while our proposed CrossQ + wn uses only $\sim 600k$ (these numbers vary slightly per environment, depending on the state and action dimensionalities).
In contrast, vanilla CrossQ utd$\mathbin{=}1$ exhibits much slower learning on most tasks and, in some environments, fails to learn performant policies. Moreover, the instability of vanilla CrossQ at utd$\mathbin{=}5$ is particularly notable, as it does not reliably converge across environments (Figure 7).
@@ 5.3 Stable scaling of CrossQ + wn with utd ratios
Figure 4: CrossQ wn utd scaling behavior. We plot the iqm return and $90\%$ confidence intervals for different utd ratios. Results are aggregated over 15 dmc environments and $10$ random seeds each according to Agarwal et al. [1]. The sample efficiency scales reliably with increasing utd ratios and remains stable even when there are no more performance gains, which is a crucial property.
To visualize the stable scaling behavior of CrossQ + wn we ablate across utd ratios. Figure 4 shows training performances aggregated over multiple environments for $10$ seeds each at 1M steps and 200k steps respectively. We confirm that CrossQ + wn shows reliable scaling behavior. With increasing compute, the performance increases or stays constant which is desirable. We see, that for the same utd ratio, CrossQ + wn nearly always beats both the bro and simba baselines.
@@ 5.4 Hyperparameter ablation studies
We also ablate the different hyperparameters of CrossQ + wn, by changing each one at a time. Figure 5 shows aggregated results of the final performances of each ablation. We will briefly discuss each ablation individually.
@@
Not performing weight normalization results in the biggest drop in performance across all our ablations. This loss is most drastic on the MyoSuite tasks and often results in no meaningful learning. Showing that, as hypothesized, the inclusion of wn into the CrossQ framework yields great improvements in terms of sample efficiency and training stability, especially for larger utd ratios. In general, lower utd ratios are already reasonably competitive in overall performance.
@@
Ablating the target networks shows that on MyoSuite, there is a small but significant difference between using a target network and or no target network. Results on dmc show a large drop in performance. There, removing target networks leads to a significant drop in performance, nearly as large as removing weight normalization. This finding is interesting, as it suggests that CrossQ + wn without target networks is not inherently unstable. But there are situations where the inclusion of target networks is required. Further investigating the role and necessity of target networks in rl is an interesting direction for future research.
L2 regularization.
Figure 5 investigates the performance of a soft L2 penalty on the weights compared to the weight normalization proposed in this paper. Across all tasks, and sweeping across soft regularization scalings, weight normalization outperforms the soft L2 penalty. Our hypothesis is that, in principle, soft L2 regularization could work in a similar way to the proposed wn; however, it would require per-task tuning and potentially even scheduling to result in a stable elr. In comparison, the hard constraint via wn guarantees a stable weight norm by design and as such is easier to employ.
Parameter resets.
Our experiments show that CrossQ + wn is able to scale without requiring drastic interventions such as parameter resets. However, it is still interesting to investigate whether CrossQ + wn’s performance could benefit from parameter resets. CrossQ + wn utd$=5$ + Reset in Figure 5 investigates this question. We see, that there is a slight improvement on the $7$ dog and humanoid tasks, on all other $18$ tasks, performance remains the same. The main takeaway is, that CrossQ + wn scales stably without requiring parameter resets.
Figure 5: An ablation study comparing CrossQ + wn against a soft L2 penality on the weights, as well as other design decisions such as target networks. The results show that the hard constraint outperforms the soft approach across a range of regularization scales and tasks. Uncertainty quantification depicts the $90\%$ stratified bootstrap confidence intervals.
@@ 6 Related work
Model-free rl methods, including those utilizing higher utd ratios, focus on increasing the number of gradient updates per collected sample to maximize learning from available data. High utd training introduces several challenges, such as overfitting to early training data, a phenomenon known as primacy bias [32]. This can be counteracted by periodically resetting the network parameters [32, 9]. However, network resets introduce abrupt performance drops. Nauman et al. [31] demonstrated that full parameter resets of the critic can effectively preserve learning capacity using a utd ratio up to $10$. However, such resets inevitably impact the wall-clock time due to relearning function approximation several times during learning. Alternative approaches use an ensemble of Q-functions to reduce overestimation bias that occurs under high utd ratio training regimes [8]. Due to the decreased computational efficiency of using a large number of Q-functions, Hiraoka et al. [17] propose to replace the ensemble of critics with dropout and layer normalization. Both methods utilize a utd ratio of 20, which is highly inefficient.
@@
Normalization techniques have long been recognized for their impact on neural network training. ln [3] and other architectural modifications have been used to stabilize learning in rl [17, 31]. Yet bn has only recently been successfully applied in this context [4], challenging previous findings, where bn in critics caused training to fail [17]. wn has been shown to keep elrs stable and prevent loss of plasticity [29], when combined with ln, making it a promising candidate for integration into existing rl frameworks.
Normalization techniques have long been recognized for their impact on neural network training.  Bjorck et al. [5] show that using spectral normalization (sn) enables training with large scale neural networks in deep rl. sn divides a layer’s weight matrix by its largest singular value, regularizing the Lipschitz continuity of the function approximation and therefore stabilizing the gradients. bn has been used by Bhatt et al. [4], where it achieves impressive results, but requires slight adjustments when the UTD ratio is scaled. Concurrent work [26] injected ‘simplicity bias’ into their actor and critic architecture, encouraging the model to favor ‘simpler’ features for its predictions. Simba incorporates a residual feedforward block and layer normalization in both the actor and critic networks and achieved state of the art results. wn has been shown to keep elrs stable and prevent loss of plasticity [29], when combined with ln, making it a promising candidate for integration into existing rl frameworks.
Hafner et al. [14] designed a vision-based mbrl algorithm Dreamer that leverages world models to master a wide range of diverse tasks, and also relies on several normalization techniques. They utilize root mean square layer normalization (RMSNorm)  [45] before the activation function and normalize the returns.
To successfully scale deep rl on the Atari 100k benchmark and achieve human-level sample efficiency,  Schwarzer et al. [36] rely on regularization techniques. Rather than use normalization methods, they use the shrink-and-perturb method [2] in shallow layers and full parameter resets in deeper layers to preserve network plasticity. To scale to higher utds, they introduce AdamW  [28] and gradually decrease the number of steps for the computation of the td error for faster convergence. Lee et al. [25] argue that the loss of plasticity observed in deep rl when the utd ratio is increased can be mitigated by using ln, sharpness-aware minimization (samForet et al. [12]), incorporating parameter resets in the last layers and replacing the ReLU activation function with a ‘concetenated ReLU’ function  [37].  Voelcker et al. [44] demonstrated that parameter resets are not strictly necessary when the utd ratio is increased to improve sample efficiency. They identify the generalization ability of the critic as the main source of training instabilities under high utd regimes. They demonstrate empirically that architectural regularization can mitigate overestimation and divergence, but it does not guarantee proper generalization. On the other hand, leveraging synthetic data generated by a learned world model can help mitigate the effects of distribution shift, thereby enhancing generalization.
7 Conclusion, limitations & future work
@@
In this work, we have addressed the instability and scalability limitations of CrossQ in rl by integrating weight normalization. Our empirical results demonstrate that wn effectively stabilizes training and allows CrossQ to scale reliably with higher utd ratios. The proposed CrossQ + wn approach achieves competitive or superior performance compared to state-of-the-art baselines across a diverse set of $25$ complex continuous control tasks from the dmc and MyoSuite benchmarks. These tasks include complex and high-dimensional humanoid and dog environments. This extension preserves simplicity while enhancing robustness and scalability by eliminating the need for drastic interventions such as network resets.
In this work, we only consider continuous state-action benchmarking tasks. While our proposed CrossQ + wn performs competitively on these tasks, its performance on discrete state-action spaces or vision-based tasks remains unexplored. We plan to investigate this in future work. Moreover, the theoretical basis of our work does not directly connect to the convergence rates or sample efficiency of the underlying rl algorithm, but rather to mitigate observed empirical phenomena regarding the function approximation alone.
Acknowledgments and Disclosure of Funding
@@
We would also like to thank Tim Schneider, Cristiana de Farias, João Carvalho and Theo Gruner for proofreading and constructive criticism on the manuscript. This research was funded by the research cluster “Third Wave of AI”, funded by the excellence program of the Hessian Ministry of Higher Education, Science, Research and the Arts, hessian.AI. This work was also supported by a UKRI/EPSRC Programme Grant [EP/V000748/1].